# Exact non-stationary solutions of the Euler equations in two and three dimensions

### v1

@@ Abstract.
We develop, via Arnold’s geometric framework, a mechanism for constructing explicit, smooth, global-in-time, and typically non-stationary solutions of the incompressible Euler equations. The approach introduces a notion of generalized Coriolis force, whose spectrum underlies the construction of these solutions. In the setting of ideal hydrodynamics, the construction recovers classical exact solutions such as Kelvin and Rossby-Haurwitz waves, while also producing new explicit examples on curved surfaces and three-dimensional manifolds including the round three-sphere. We further obtain a complete classification in two dimensions and a partial classification in three dimensions of the Riemannian manifolds that admit such solutions. The method is also formulated in the general Euler-Arnold setting and yields a simple criterion for non-stationarity.
@@ 1. Introduction
(1.1) $\begin{split}&\partial_{t}u+\nabla_{u}u=-\nabla p,\\ &\mathrm{div}u=0,\end{split}$
@@
Arnold’s 1966 work [3] recast ideal fluid motion in geometric terms. In this formulation, the Euler equations coincide with the reduced geodesic equation on the group of volume-preserving diffeomorphisms equipped with a right-invariant $L^{2}$ metric arising from kinetic energy. Ebin and Marsden [15] proved that the corresponding geodesic flow in the $H^{s}$ Sobolev setting is governed by an ordinary differential equation and obtained smoothness of the associated Riemannian exponential map, which may be identified with the Lagrangian flow map of (1.1). This geometric formulation has since played a central role in the study of stability, curvature, and qualitative features of fluid motion. An extensive treatment of this geometric approach can be found in the monograph of Arnold and Khesin [2].
@@
The central result of this paper utilizes Arnold’s geometric framework to provide a general mechanism for constructing a class of explicit, smooth, global-in-time and typically non-stationary solutions to the Euler equations (1.1) and, more broadly, to Euler-Arnold equations on Lie groups. We also give a straightforward criterion guaranteeing non-stationarity of the solutions. The underlying mechanism can be understood from two complementary viewpoints.
The first viewpoint is perturbative, though no smallness assumption is imposed on the size of the perturbation. One begins with a Killing field $X$ whose flow, by definition, acts by isometries and hence produces a steady Euler flow. Introducing a perturbation of this steady solution and formally expanding the resulting solution of (1.1) leads to an infinite system of coupled ordinary differential equations governing the evolution of the expansion coefficients. A particularly simple class of solutions to this system arises when the perturbation is chosen from a complex simultaneous eigenfield of the inertia operator [^1] For ideal hydrodynamics, this is the Hodge Laplacian when the fluid is two-dimensional and the curl operator when the fluid is three-dimensional and the Lie algebra coadjoint operator associated with $X$. In this situation the nonlinear interactions close on a finite-dimensional subspace, producing explicit time-dependent solutions. Importantly, this construction yields exact solutions of the fully nonlinear Euler equations (1.1), rather than solutions of a linearized model.
@@
When applied to two-dimensional ideal hydrodynamics our construction recovers several well-known solutions, including Kelvin waves on flat domains and Rossby-Haurwitz waves on the round sphere. In addition, we obtain explicit solutions on curved surfaces such as the hyperbolic disk. In three dimensions we produce new explicit solutions of the Euler equations on manifolds including the round three-sphere and various circle and torus bundles with nontrivial geometry. On the three-sphere, these solutions may be viewed as higher-dimensional analogues of Rossby-Haurwitz waves: they are periodic in time and occur in families of multiplicity greater than one.
@@
A particularly transparent version of our algorithm arises when the chosen Killing field has the additional property that its image under the inertia operator is also Killing. For example, on a flat cylinder, the curl of the rotational vector field generates vertical translations, which are themselves isometries. All explicit examples presented in this paper fall into this category. We therefore undertake a systematic study of manifolds admitting such Killing fields. In two dimensions, we obtain a complete classification, while in three dimensions we give a partial classification.
@@
The paper is organized as follows. Section 2 introduces the general mechanism for constructing solutions in the context of ideal hydrodynamics and formulates sufficient conditions under which the underlying manifold admits an abundance of such solutions. Sections 3 and 4 provide a full and partial classification, respectively, of two- and three-dimensional manifolds satisfying these conditions, together with explicit examples. Section 5 develops the construction in the general Euler-Arnold setting. Lastly, Section 6 outlines directions for future work.
@@ Acknowledgments
The authors thank Abhijit Champanerkar, Theodore Drivas, Gerard Misiołek, and Daniel Peralta-Salas for helpful discussions.
@@ 2. Ideal Hydrodynamics
Let $(M,g)$ be a two- or three-dimensional compact Riemannian manifold, possibly with boundary. Denote by $\mathfrak{X}_{\mu}(M)$ the space of smooth divergence-free vector fields on $M$ that are tangent to the boundary. Within $\mathfrak{X}_{\mu}(M)$ there is the finite-dimensional subspace of harmonic fields $\mathcal{H}(M)$, consisting of divergence-free and curl-free fields tangent to the boundary. By the Hodge decomposition (cf. [15]) we have the $L^{2}$-orthogonal splitting
@@
where $\mathfrak{X}_{\mu,\text{ex}}(M)$ denotes the $L^{2}$-orthogonal complement of $\mathcal{H}(M)$ in $\mathfrak{X}_{\mu}(M)$. We denote the usual musical isomorphisms $\flat$ and $\sharp$, the Hodge star operator $\star$, the exterior derivative $d$, and its formal adjoint $\delta$.
@@
where $\Delta$ is the Hodge Laplacian
$\Delta u=(d\delta+\delta d)(u^{\flat})^{\sharp},$
@@
$\operatorname{curl}u=(\star(du^{\flat}))^{\sharp}.$
@@
Let $(M,g)$ be a two- or three-dimensional compact Riemannian manifold, possibly with boundary. The inertia operator $\mathcal{A}$ given in (2.1) is invertible on $\mathfrak{X}_{\mu,\text{ex}}(M)$ and its inverse $\mathcal{A}^{-1}$ is compact and self-adjoint in $L^{2}$. Consequently, $\mathcal{A}$ has a discrete spectrum $\{\alpha_{k}\}$ with finite-dimensional eigenspaces, and the corresponding eigenfields form an $L^{2}$-orthonormal basis of $\mathfrak{X}_{\mu,\text{ex}}(M)$.
@@
In two dimensions, as noted above, for each $v\in\mathfrak{X}_{\mu,\text{ex}}$, we have $v=\nabla^{\perp}\psi$, for some stream function $\psi$. Hence, we have $\mathcal{A}\nabla^{\perp}\phi=\nabla^{\perp}\Delta\phi,$ so that $\mathcal{A}^{-1}(\nabla^{\perp}\phi)=\nabla^{\perp}\Delta_{0}^{-1}\phi$, where $\Delta_{0}^{-1}$ denotes the inverse Laplacian with Dirichlet boundary conditions on the stream function. The $L^{2}$ norm on skew-gradients coincides with the Dirichlet norm on the stream functions, and $\Delta_{0}^{-1}$ is compact and self-adjoint in this inner product.
@@
where for $v,w\in\mathfrak{X}_{\mu,\text{ex}}(M)$ the bracket [v,w]=vwwv[v,w]=vw-wv denotes the standard commutator of vector fields. In accordance with the framework outlined in Section 5, we will refer to the operators $\operatorname{ad}_{u_{0}}:v\mapsto-[u_{0},v]$ and $v\mapsto K_{u_{0}}v$ as the adjoint and coadjoint operators associated to $u_{0}$ respectively. The operator $K_{u_{0}}$ is compact if the fluid is two-dimensional while it is generally only bounded if the fluid is three-dimensional, see [16] for further discussion and consequences.
With this notation, the Euler equations (1.1) may be written in the form
(2.3) $\partial_{t}\mathcal{A}u+[u,\mathcal{A}u]=0,\qquad\mathrm{div}u=0,\qquad u\parallel\partial M.$
@@
Let $(M,g)$ be a Riemannian manifold, possibly with boundary, and let $u_{0}\in\mathfrak{X}_{\mu,\text{ex}}(M)$ be a steady solution of the Euler equations (1.1), i.e. $u_{0},\mathcal{A}u_{0}]=0$. If there exists a complex vector field $z=v+iw\in\mathbb{C}\otimes\mathfrak{X}_{\mu,\text{ex}}(M)$, with $v,w\in\mathfrak{X}_{\mu,\text{ex}}(M)$, and real constants $\lambda,\alpha,\zeta$ such that
$K_{u_{0}}z=-i\lambda z,\qquad\mathcal{A}z=\alpha z,\qquad[u_{0},z]=i\zeta z,$
then the curve
$U(t)=u_{0}+e^{i(\lambda-\zeta)t}z$
@@
$U_{\mathfrak{R}}(t)=u_{0}+\mathrm{Re}\big(e^{i(\lambda-\zeta)t}z\big)=u_{0}+\cos\big((\lambda-\zeta)t\big)v-\sin\big((\lambda-\zeta)t\big)w$
@@
$U_{\mathfrak{I}}(t)=\mathrm{Im}\big(e^{i(\lambda-\zeta)t}z\big)=\sin\big((\lambda-\zeta)t\big)v+\cos\big((\lambda-\zeta)t\big)w$
@@
In fact, for each such field $z$ we have a two-dimensional space of such solutions which corresponds to the complex span of $z$. Replacing $z$ with $\rho e^{i\sigma}z$ for $\rho\geq 0$ and $\sigma\in\mathbb{R}$ yields the entire set of solutions generated by $z$
$U(t)$ $=u_{0}+\rho e^{i\big(\sigma+(\lambda-\zeta)t\big)}z$
$U_{\mathfrak{R}}(t)$ $=u_{0}+\mathrm{Re}\Big(\rho e^{i\big(\sigma+(\lambda-\zeta)t\big)}z\Big)=u_{0}+\rho\cos\big(\sigma+(\lambda-\zeta)t\big)v-\rho\sin\big(\sigma+(\lambda-\zeta)t\big)w$
$U_{\mathfrak{I}}(t)$ $=\mathrm{Im}\Big(\rho e^{i\big(\sigma+(\lambda-\zeta)t\big)}z\Big)=\rho\sin\big(\sigma+(\lambda-\zeta)t\big)v+\rho\cos\big(\sigma+(\lambda-\zeta)t\big)w.$
However, we will continue in this section, and indeed in Sections 3 and 4, with $\rho=1$ and $\sigma=0$ for simplicity.
@@
$\partial_{t}\mathcal{A}U=\partial_{t}\left(\mathcal{A}u_{0}+\alpha e^{i(\lambda-\zeta)t}z\right)=i\alpha(\lambda-\zeta)e^{i(\lambda-\zeta)t}z$
@@
$[U,\mathcal{A}U]$ $=\left[u_{0}+e^{i(\lambda-\zeta)t}z,\mathcal{A}u_{0}+\alpha e^{i(\lambda-\zeta)t}z\right]$
$=\left[u_{0},\mathcal{A}u_{0}\right]+e^{i(\lambda-\zeta)t}\Big(\alpha\left[u_{0},z\right]+\left[z,\mathcal{A}u_{0}\right]\Big)+e^{2i(\lambda-\zeta)t}[z,z]$
$=e^{i(\lambda-\zeta)t}\Big(i\alpha\zeta z+\mathcal{A}K_{u_{0}}z\Big)$
$=i\alpha(\zeta-\lambda)e^{i(\lambda-\zeta)t}z$
and hence $U$ is a complex-valued solution of the Euler equations (1.1).
@@
Notice now that $\mathcal{A}z=\alpha z$ implies both $\mathcal{A}v=\alpha v$ and $\mathcal{A}w=\alpha w$ which gives us $\mathcal{A}U_{\mathfrak{I}}=\alpha U_{\mathfrak{I}}$. Hence $U_{\mathfrak{I}},\mathcal{A}U_{\mathfrak{I}}]=0$ and we see that $U_{\mathfrak{R}}$ is a real solution of the Euler equations (1.1). Meanwhile the second equation is precisely the statement that $U_{\mathfrak{I}}$ satisfies the linearized Euler equations at $U_{\mathfrak{R}}$.
@@
As will be explained in more detail in Section 5, the flow of our solution $U_{\mathfrak{R}}$ and the flow of the field $V(t)=\mathrm{Re}\left(e^{i\lambda t}z\right)$ are related by the equation
$\gamma_{U}=\gamma_{X}\circ\gamma_{V}.$
Notice that it may occur that $U_{\mathfrak{R}}$ is a non-stationary solution of (1.1), while also having that $V$ is a stationary field. In this situation $U_{\mathfrak{R}}$ can be interpreted as a trivial solution in a moving frame generated by $X$; like coffee being stirred on a moving train. However, we see that
@@
Accordingly, eigenfields $z$ with $\lambda=0$ produce solutions that are trivial in this moving-frame sense, whereas $\lambda\neq 0$ yields solutions $U_{\mathfrak{R}}$ that are genuinely non-trivial. This corresponds to the importance of vorticity twisting.
We now turn to the abundance of such solutions. Let $X\in\mathfrak{X}_{\mu,\text{ex}}(M)$ be a Killing field on $(M,g)$, that is, a vector field whose flow $\gamma_{X}$ acts by isometries. Any Killing field is automatically a steady solution of (1.1), cf. [24]. We show that if both $u_{0}$ and $\mathcal{A}u_{0}$ are Killing, then there exists an infinite-dimensional family of $z\in\mathbb{C}\otimes\mathfrak{X}_{\mu,\text{ex}}(M)$ satisfying the hypotheses of Theorem 2.2, and hence an infinite collection of exact, non-stationary solutions of (1.1). We begin with a structural lemma.
Lemma 2.5.
(2.4) $\langle\mathcal{A}^{-1}u,[v,w]\rangle_{L^{2}}+\langle\mathcal{A}^{-1}w,[v,u]\rangle_{L^{2}}=0.$

### v2

@@ Abstract.
We develop, via Arnold’s geometric framework, a mechanism for constructing explicit, smooth, global-in-time, and typically non-stationary solutions of the incompressible Euler equations. The approach introduces a notion of generalized Coriolis force, whose spectrum underlies the construction of these solutions. We recover classical exact solutions such as Kelvin and Rossby-Haurwitz waves, while also producing new explicit examples on curved surfaces and three-dimensional manifolds including the round three-sphere. Furthermore, we obtain a complete classification in two dimensions and a partial classification in three dimensions of the Riemannian manifolds that admit such solutions. The method is in fact formulated in the general Euler-Arnold setting and yields a simple criterion for non-stationarity.
@@ 1. Introduction
(1.1) $\begin{split}&\partial_{t}u+\nabla_{u}u=-\nabla p,\\ &\operatorname{div}u=0,\end{split}$
@@
Arnold’s 1966 work [3] recast ideal fluid motion in geometric terms. In this formulation, the Euler equations coincide with the reduced geodesic equation on the group of volume-preserving diffeomorphisms equipped with a right-invariant $L^{2}$ metric arising from kinetic energy. Ebin and Marsden [17] proved that the corresponding geodesic flow in the $H^{s}$ Sobolev setting is governed by an ordinary differential equation and obtained smoothness of the associated Riemannian exponential map, which may be identified with the Lagrangian flow map of (1.1). This geometric formulation has since played a central role in the study of stability, curvature, and qualitative features of fluid motion. An extensive treatment of this approach can be found in the monograph of Arnold and Khesin [2].
@@
The central result of this paper utilizes Arnold’s geometric framework to provide a general mechanism for constructing a class of explicit, smooth, global-in-time and typically non-stationary solutions to the Euler equations (1.1) and, more broadly, to Euler-Arnold equations on Lie groups. The underlying mechanism can be understood from two complementary viewpoints.
The first viewpoint is perturbative, though no smallness assumption is imposed on the size of the perturbation. One begins with a Killing field $X$ whose flow, by definition, acts by isometries and hence produces a steady Euler flow. Introducing a perturbation of this steady solution and formally expanding the resulting solution of (1.1) leads to an infinite system of coupled ordinary differential equations governing the evolution of the expansion coefficients. We demonstrate that if the second-order term vanishes, all higher-order terms vanish. Hence the solution of the linearized Euler equation is also a solution of the full Euler equation.
A particularly simple class of solutions to this system arises when the perturbation is chosen from a complex simultaneous eigenfield of the inertia operator [^1] For ideal hydrodynamics, this is the Hodge Laplacian when the fluid is two-dimensional and the curl operator when the fluid is three-dimensional. and the Lie algebra coadjoint operator associated with $X$. In this situation the nonlinear interactions close on a finite-dimensional subspace, producing explicit time-dependent solutions. Importantly, this construction yields exact solutions of the fully nonlinear Euler equations (1.1), rather than solutions of a linearized model.
@@
When applied to two-dimensional ideal hydrodynamics our construction recovers several well-known solutions, including Kelvin waves on flat domains and Rossby-Haurwitz waves on the round sphere [^2] Note that some of these have been independently derived as Euler solutions by other authors [4, 11]. In addition, we obtain explicit solutions on curved surfaces such as the hyperbolic disk. In three dimensions we produce new explicit solutions of the Euler equations on manifolds including the round three-sphere and various circle and torus bundles with nontrivial geometry. On the three-sphere, these solutions may be viewed as higher-dimensional analogues of Rossby-Haurwitz waves: they are periodic in time and occur in families of multiplicity greater than one.
@@
In all of our examples the chosen Killing field has the additional property that its image under the inertia operator is also Killing. For example, on a flat cylinder, the curl of the rotational vector field generates vertical translations, which are themselves isometries. All explicit examples presented in this paper fall into this category. We therefore undertake a systematic study of manifolds admitting such Killing fields. In two dimensions, we obtain a complete classification, while in three dimensions we give a partial classification.
@@
The paper is organized as follows. Section 2 introduces the general mechanism for constructing solutions in the context of ideal hydrodynamics and formulates sufficient conditions under which the underlying manifold admits an abundance of such solutions. Sections 3 and 4 provide a full and partial classification, respectively, of two- and three-dimensional manifolds satisfying these conditions, together with explicit examples. Section 5 develops the construction in the general Euler-Arnold setting. Section 6 outlines directions for future work. Lastly, Appendix A contains proofs of a Riemannian geometric nature which are deferred for readability.
@@ Acknowledgments
The authors thank Abhijit Champanerkar, Albert Chern, Theodore Drivas, Boris Khesin, Gerard Misiołek, Daniel Peralta-Salas and Cornelia Vizman for helpful discussions and comments.
@@ 2. Ideal Hydrodynamics
Let $(M,g)$ be a two- or three-dimensional compact Riemannian manifold, possibly with boundary. Denote by $\mathfrak{X}_{\mu}(M)$ the space of smooth divergence-free vector fields on $M$ that are tangent to the boundary. Within $\mathfrak{X}_{\mu}(M)$ there is the finite-dimensional subspace of harmonic fields $\mathcal{H}(M)$, consisting of divergence-free and curl-free fields tangent to the boundary. By the Hodge decomposition (cf. [17]) we have the splitting
@@
where $\mathfrak{X}_{\mu,\text{ex}}(M)$ denotes the $L^{2}$-orthogonal complement of $\mathcal{H}(M)$ in $\mathfrak{X}_{\mu}(M)$. We recall the usual musical isomorphisms $\flat$ and $\sharp$, the Hodge star operator $\star$, the exterior derivative $d$, and its formal adjoint $\delta$.
@@
where $\Delta$ is the positive-definite Hodge Laplacian
(2.2) $\Delta u=(d\delta+\delta d)(u^{\flat})^{\sharp},$
@@
(2.3) $\operatorname{curl}u=(\star(du^{\flat}))^{\sharp}.$
@@
Let $(M,g)$ be a two- or three-dimensional compact Riemannian manifold, possibly with boundary. The inertia operator $\mathcal{A}$ given in (2.1) is invertible on $\mathfrak{X}_{\mu,\text{ex}}(M)$, and its inverse $\mathcal{A}^{-1}$ is compact and self-adjoint in $L^{2}$. Consequently, $\mathcal{A}$ has a discrete spectrum $\{\alpha_{k}\}$ with finite-dimensional eigenspaces, and the corresponding eigenfields form an $L^{2}$-orthonormal basis of $\mathfrak{X}_{\mu,\text{ex}}(M)$.
@@
In two dimensions, for each $v\in\mathfrak{X}_{\mu,\text{ex}}$, we have $v=\nabla^{\perp}\psi$, for some stream function $\psi$ vanishing on all boundary components. Hence, we have $\mathcal{A}\nabla^{\perp}\psi=\nabla^{\perp}\Delta\psi,$ so that $\mathcal{A}^{-1}(\nabla^{\perp}\psi)=\nabla^{\perp}\Delta^{-1}\psi$, where $\Delta^{-1}$ denotes the inverse Laplacian with Dirichlet boundary conditions on the stream function. The $L^{2}$ norm on skew-gradients coincides with the Dirichlet norm on the stream functions, and $\Delta^{-1}$ is compact and self-adjoint in this inner product, cf. [19].
@@
where for $v,w\in\mathfrak{X}_{\mu,\text{ex}}(M)$ the bracket [v,w]=vwwv[v,w]=vw-wv denotes the standard commutator of vector fields. In accordance with the framework outlined in Section 5, we will refer to the operators $\operatorname{ad}_{u_{0}}:v\mapsto-[u_{0},v]$ and $v\mapsto K_{u_{0}}v$ as the adjoint and coadjoint operators associated to $u_{0}$ respectively. The operator $K_{u_{0}}$ is compact if the fluid is two-dimensional, while it is generally only bounded if the fluid is three-dimensional; see [18] for further discussion and consequences.
With this notation, the Euler equations (1.1) can be written in the form
(2.5) $\partial_{t}\mathcal{A}u+[u,\mathcal{A}u]=0,\qquad\operatorname{div}u=0,\qquad u\parallel\partial M.$
@@
Let $(M,g)$ be a Riemannian manifold, possibly with boundary, and let $u_{0}\in\mathfrak{X}_{\mu,\text{ex}}(M)$ be a steady solution of the Euler equations (1.1), i.e., $u_{0},\mathcal{A}u_{0}]=0$. If there exists a complex vector field $\mathfrak{q}=v+iw\in\mathbb{C}\otimes\mathfrak{X}_{\mu,\text{ex}}(M)$, with $v,w\in\mathfrak{X}_{\mu,\text{ex}}(M)$, and real constants $\lambda,\alpha,\zeta$ such that
$K_{u_{0}}\mathfrak{q}=-i\lambda\mathfrak{q},\qquad\mathcal{A}\mathfrak{q}=\alpha\mathfrak{q},\qquad[u_{0},\mathfrak{q}]=i\zeta\mathfrak{q},$
then, defining $\varpi=\lambda-\zeta$, the curve
$U_{\mathfrak{C}}(t)=u_{0}+e^{i\varpi t}\mathfrak{q}$
@@
$U_{\mathfrak{R}}(t)=u_{0}+\mathrm{Re}\big(e^{i\varpi t}\mathfrak{q}\big)=u_{0}+\cos(\varpi t)v-\sin(\varpi t)w$
@@
$U_{\mathfrak{I}}(t)=\mathrm{Im}\big(e^{i\varpi t}\mathfrak{q}\big)=\sin(\varpi t)v+\cos(\varpi t)w$
@@
In fact, for each field $\mathfrak{q}$ we have a two-dimensional space of such solutions which corresponds to the complex span of $\mathfrak{q}$. Replacing $\mathfrak{q}$ with $\rho e^{i\sigma}\mathfrak{q}$ for $\rho\geq 0$ and $\sigma\in\mathbb{R}$ yields the entire set of solutions generated by $\mathfrak{q}$
$U_{\mathfrak{C}}(t)$ $=u_{0}+\rho e^{i(\sigma+\varpi t)}\mathfrak{q}$
$U_{\mathfrak{R}}(t)$ $=u_{0}+\mathrm{Re}\Big(\rho e^{i(\sigma+\varpi t)}\mathfrak{q}\Big)=u_{0}+\rho\cos(\sigma+\varpi t)v-\rho\sin(\sigma+\varpi t)w$
$U_{\mathfrak{I}}(t)$ $=\mathrm{Im}\Big(\rho e^{i(\sigma+\varpi t)}\mathfrak{q}\Big)=\rho\sin(\sigma+\varpi t)v+\rho\cos(\sigma+\varpi t)w.$
The parameter $\rho$ governs the size of the perturbation, relative to the fixed steady flow $u_{0}$, while $\sigma$ gives a phase. However, we will continue in this section, and indeed largely in Sections 3 and 4, with $\rho=1$ and $\sigma=0$ for simplicity.
@@
$\partial_{t}\mathcal{A}U_{\mathfrak{C}}=\partial_{t}\left(\mathcal{A}u_{0}+\alpha e^{i\varpi t}\mathfrak{q}\right)=i\alpha\varpi e^{i\varpi t}\mathfrak{q}$
@@
$[U_{\mathfrak{C}},\mathcal{A}U_{\mathfrak{C}}]$ $=\left[u_{0}+e^{i\varpi t}\mathfrak{q},\mathcal{A}u_{0}+\alpha e^{i\varpi t}\mathfrak{q}\right]$
$=\left[u_{0},\mathcal{A}u_{0}\right]+e^{i\varpi t}\Big(\alpha\left[u_{0},\mathfrak{q}\right]+\left[\mathfrak{q},\mathcal{A}u_{0}\right]\Big)+e^{2i\varpi t}[\mathfrak{q},\mathfrak{q}]$
$=e^{i\varpi t}\Big(i\alpha\zeta\mathfrak{q}+\mathcal{A}K_{u_{0}}\mathfrak{q}\Big)$
$=i\alpha(\zeta-\lambda)e^{i\varpi t}\mathfrak{q}$
$=-i\alpha\varpi e^{i\varpi t}\mathfrak{q}$
and hence $U_{\mathfrak{C}}$ is a complex-valued solution of the Euler equations (1.1).
@@
Notice now that $\mathcal{A}\mathfrak{q}=\alpha\mathfrak{q}$ implies both $\mathcal{A}v=\alpha v$ and $\mathcal{A}w=\alpha w$ which gives us $\mathcal{A}U_{\mathfrak{I}}=\alpha U_{\mathfrak{I}}$. Hence $U_{\mathfrak{I}},\mathcal{A}U_{\mathfrak{I}}]=0$ and we see that $U_{\mathfrak{R}}$ is a real solution of the Euler equations (1.1). Meanwhile the second equation is precisely the statement that $U_{\mathfrak{I}}$ satisfies the linearized Euler equations at $U_{\mathfrak{R}}$.
@@
As will be explained in more detail in Section 5, the flow of our solution $U_{\mathfrak{R}}$ and the flow of the field $V(t)=\mathrm{Re}\left(e^{i\lambda t}\mathfrak{q}\right)$ are related by the equation
$\gamma_{U_{\mathfrak{R}}}=\gamma_{u_{0}}\circ\gamma_{V}.$
Notice that it may occur that $U_{\mathfrak{R}}$ is a non-stationary solution of (1.1), while also having that $V$ is a stationary field. In this situation $U_{\mathfrak{R}}$ can be interpreted as a trivial solution in a moving frame generated by $u_{0}$; like coffee being stirred on a moving train. However, we see that
@@
Accordingly, eigenfields $\mathfrak{q}$ with $\lambda=0$ produce solutions that are trivial in this moving-frame sense, whereas $\lambda\neq 0$ yields solutions $U_{\mathfrak{R}}$ that are genuinely non-trivial. This corresponds to the importance of vorticity twisting.
We now turn to the abundance of such solutions. A vector field $X\in\mathfrak{X}_{\mu,\text{ex}}(M)$ is said to be a Killing field if its flow $\gamma_{X}$ acts by isometries. Equivalently, for all $u,v\in\mathfrak{X}_{\mu,\text{ex}}(M)$ we have
(2.6) $g(\nabla_{u}X,v)+g(u,\nabla_{v}X)=0.$
As we will demonstrate in the appendix, any such field is automatically a steady solution of (1.1), cf. [27].
Remark 2.5.
If the vector field $u_{0}$ in Theorem 2.2 is in fact Killing and the complex field $\mathfrak{q}$ is such that $\zeta=0$, then the corresponding solution $U_{\mathfrak{R}}$ is axi-symmetric with respect to $u_{0}$.
Furthermore, for simplicity in the spectral theory, we have restricted to $\mathfrak{q}\in\mathbb{C}\otimes\mathfrak{X}_{\mu,\text{ex}}(M)$, the complexified $L^{2}$-orthogonal complement of harmonic fields within divergence-free fields; hence we we will not have $\alpha=0$ for non-trivial $\mathfrak{q}$. However we note that there exist exact solutions of (1.1) such that $\mathcal{A}u$ is steady, but $u$ has an oscillating harmonic component, cf. Yin et al. [41]. Such dynamics may be interesting in this context.
Our central ingredient is the following.
Proposition 2.6.
Let $(M,g)$ be a compact Riemannian manifold, possibly with boundary, and $X\in\mathfrak{X}_{\mu,\text{ex}}(M)$ be a Killing field such that $\mathcal{A}X$ is also Killing. Then the operators
$K_{X}:\mathfrak{q}\mapsto K_{X}\mathfrak{q},\quad\mathcal{A}:\mathfrak{q}\mapsto\mathcal{A}\mathfrak{q},\quad\operatorname{ad}_{X}:\mathfrak{q}\mapsto-[X,\mathfrak{q}]$
In particular, if both $X$ and $\mathcal{A}X$ are Killing, then there exists an infinite-dimensional family of $\mathfrak{q}\in\mathbb{C}\otimes\mathfrak{X}_{\mu,\text{ex}}(M)$ satisfying the hypotheses of Theorem 2.2, with $u_{0}=X$, and hence an infinite collection of exact, non-stationary solutions of (1.1). We present the proof in Appendix A.
Theorem 2.7.
Let $(M,g)$ be a compact Riemannian manifold, possibly with boundary and denote by $\mathfrak{X}_{\mu,\text{ex}}(M)$ those divergence-free vector fields which are tangent to the boundary of $M$ and $L^{2}$-orthogonal to harmonic fields.
If there exists a Killing field $X\in\mathfrak{X}_{\mu,\text{ex}}(M)$ such that $\mathcal{A}X$ is also Killing, where
$\mathcal{A}=\begin{cases}\Delta&\dim(M)=2,\\ \operatorname{curl}&\dim(M)=3,\end{cases}$
then there exists infinitely many $v,w\in\mathfrak{X}_{\mu,\text{ex}}(M)$ with constants $\lambda,\alpha$ and $\zeta$ such that
$\mathcal{A}^{-1}[v,\mathcal{A}X]$ $=\lambda w,$ $\quad\mathcal{A}v$ $=\alpha v,$ $\quad[X,v]$ $=-\zeta w,$
$\mathcal{A}^{-1}[w,\mathcal{A}X]$ $=-\lambda v,$ $\quad\mathcal{A}w$ $=\alpha w,$ $\quad[X,w]$ $=\zeta v.$
Consequently, for all $\rho\geq 0$ and $\sigma\in\mathbb{R}$, we have a real solution of the Euler equations (1.1) given by
(2.7) $U_{\mathfrak{R}}(t)=X+\rho\cos(\sigma+\varpi t)v-\rho\sin(\sigma+\varpi t)w$
where $\varpi=\lambda-\zeta$.